In a science lab, 5.00 g of NaOH(s) is dissolved in 300 mL of water. What is the molar concentration of 

the resulting solution?


1
Expert's answer
2021-11-29T04:38:55-0500

To get molar concentration we use the formula:

n=CV

Where n is number of mole

C is molar concentration

V is volume

C=n/V

To get n, n=mass/molar mass

n=5/40

n=0.125

C=0.125/0.3

C=0.417M